### v1

# Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in

Anatomy of singlet-doublet dark matter relic: annihilation, co-annihilation, co-scattering, and freeze-in
[…]
${}^{1}\,$ Department of Physics, Indian Institute of Technology Hyderabad, Kandi, Sangareddy 502285, Telangana, India
The singlet-doublet vector-like fermion dark matter model has been extensively studied in the literature over the past decade. An important parameter in this model is the singlet-doublet mixing angle ($\sin\theta$). All the previous studies have primarily focused on annihilation and co-annihilation processes for obtaining the correct dark matter relic density, assuming that the singlet and doublet components decouple at the same epoch. In this work, we demonstrate that this assumption holds only for larger mixing angles with a dependency on the mass of the dark matter. However, it badly fails for the mixing angle $\sin\theta<0.05$. We present a systematic study of the parameter space of the singlet-doublet dark matter relic, incorporating annihilation, co-annihilation, and, for the first time, co-scattering processes. Additionally, the freeze-in parameter space is also explored. We found that due to the inclusion of co-scattering processes, the correct relic density parameter space is shifted towards the detection sensitivity range of the LHC and MATHUSLA via displaced vertex signatures.
1 Introduction
Dark matter (DM) is the mysterious form of matter that makes up about 26.8% of the Universe’s total energy density [1, 2]. Its existence on large scales has been firmly established through astrophysical evidence such as galaxy rotation curves, gravitational lensing, and the large-scale structure of the Universe. As the standard model (SM) of particle physics is unable to provide a suitable candidate for DM, particle physicists look beyond the SM (BSM) scenarios for DM. Weakly Interacting Massive Particle (WIMP) [3, 4, 5] is one of the well-motivated BSM scenarios where the DM relic density is established due to the freeze-out of different processes in the early Universe.
A simplest extension of the SM for a DM candidate is a vector-like singlet fermion: $\chi$ with an additional $\mathcal{Z}_{2}$ symmetry under which $\chi$ is odd while all other particles are even [8]. It has only a dimension five effective coupling with the SM Higgs boson , i.e. $(\bar{\chi}\chi H^{\dagger}H)/\Lambda$, where $\Lambda$ is the cutoff scale. A very large coupling is required to satisfy the correct relic density. However, the direct detection experiments exclude such large couplings. In the singlet fermionic DM model, only DM mass at the Higgs resonance ($M_{\rm DM}=M_{H}/2$) survives the direct detection constraints. On the other hand, if the SM is extended with a $\mathcal{Z}_{2}$-odd vector-like fermion doublet, then the neutral component of the doublet can be a viable DM candidate[8]. However, due to large gauge coupling, the annihilation cross-section is so large that the DM becomes under-abundant in the low mass region. The relic is satisfied at a very high DM mass, typically at the TeV scale. Moreover, the doublet DM model is ruled out by the current direct detection constraints [6]. This problem can be resolved if the singlet is combined with the vector-like doublet fermion, which are odd under a $\mathcal{Z}_{2}$ symmetry. In this case, the neutral component of the doublet mixes with the singlet to give a singlet-doublet DM (SDDM). By choosing the SD mixing appropriately, one can get the correct relic density in a large range of parameter space, which also satisfies the direct detection constraints. The SDDM has been studied extensively in the literature [7, 9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. An interesting feature of this model is that it introduces co-annihilation processes along with the annihilation among the dark sector particles to give rise the correct relic density of DM. The two most important parameters in the SDDM model are the SD mixing angle ($\sin\theta$) and the mass splitting ($\Delta{M}$) between the DM (dominantly singlet) and the NLSP (dominantly the doublet). For a larger $\sin\theta$ and $\Delta M$, the annihilation process dominantly decides the relic density of the DM. On the other hand, in the small mass-splitting regime, the co-annihilation process becomes dominant. In previous studies of the SDDM [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18], it was generally assumed that the dark sector particles decouple from the thermal bath at the same epoch, independent of $\sin\theta$. However, this assumption may not hold true for small mixing angles; typically, for $\sin\theta\lesssim 0.05$, the DM particles are not in chemical equilibrium with the thermal bath. In particular, the singlet component of the dark sector decouples from the thermal bath at an earlier time, while the doublet component remains in equilibrium for a longer duration due to its interactions with the SM gauge bosons. This implies that for $\sin\theta\lesssim 0.05$, the relic density parameter space estimated in earlier studies is not valid, while they remain valid for $\sin\theta\gtrsim 0.05$
In this paper, we revisit the SDDM relic density in the small mixing limits, typically $10^{-6}\lesssim\sin\theta\lesssim 0.05$. In this range of SD mixing, the singlet can be brought to equilibrium quickly. However, the DM (dominantly singlet) decouples early with a larger abundance. We demonstrate that by incorporating co-scattering or conversion-driven processes [32, 33], the DM relic density can be brought into agreement with the observed value. In particular, the kinetic equilibrium between DM and the thermal bath allows for upscattering processes, where DM particles can be scattered into a higher-mass state within the dark sector, such as the doublet, thereby reducing the singlet’s abundance. Recent studies have explored co-scattering processes in various BSM contexts; see, for instance, [34, 35, 36].
We further explore the DM parameter space by reducing the SD mixing angle. We saw that for $\sin\theta<\mathcal{O}(10^{-6})$, the singlet never reaches equilibrium, and hence the DM relic density can not be realized via the freeze-out mechanism. In such a small mixing angle regime, the DM can be produced non-thermally from the decay of the doublet partner. In Section 3.3, we saw that for DM mass ranging from 1 to 1000 GeV, the required $\sin\theta$ for correct relic density varies from $10^{-13}$ to $10^{-7}$.
We demonstrate that in the small mixing limit, the SDDM can yield distinctive signatures at colliders. The doublet fermion can be produced copiously due to their gauge interaction. Once produced, they can decay to DM and leptons via the singlet doublet mixing, which can give displaced vertex signatures in the LHC [37] and, in future, MATHUSLA [38].
The paper is organized as follows. We discuss the SDDM model in Section 2. In Section 3.1, we study the correct relic parameter space considering annihilation and co-annihilation only. We study the co-scattering effects in Section 3.2. The freeze-in production of DM is discussed in Section 3.3. The constraint on the DM parameter space using the null detection of the DM at direct search experiments is discussed in Section 4. Section 5 discusses the displaced vertex signatures. We finally conclude in Section 6.
2 The singlet-doublet dark matter model
In this case, the SM is extended with a fermion doublet $\Psi=(\psi^{0}~{}~{}\psi^{-})^{T}$ and a singlet fermion $\chi$, which are stabilized by a discrete symmetry $\mathcal{Z}_{2}$, under which they are odd, and all other SM particles are charged even [8].
The relevant Lagrangian of the model guided by imposed symmetry is given by
$\mathcal{L}$ $\supset$ $i\overline{\Psi}\gamma^{\mu}D_{\mu}\Psi+i\overline{\chi}\gamma^{\mu}\partial_{\mu}\chi-M_{\Psi}\overline{\Psi}\Psi-M_{\chi}\overline{\chi}\chi-y\overline{\Psi}\tilde{H}\chi+h.c.,$ (1)
where $D_{\mu}=\partial_{\mu}-g\frac{\tau_{i}}{2}W^{i}_{\mu}-g^{\prime}\frac{Y}{2}B_{\mu}$ and $H$ is the SM Higgs doublet.
The scalar potential of the model is given by
$V_{\rm scalar}$ $=$ $-\mu^{2}_{H}(H^{\dagger}H)+\lambda_{H}(H^{\dagger}H)^{2}.$ (2)
After the electro-weak symmetry breaking, the quantum fluctuation around the minima is given as
$H=\begin{pmatrix}0\\ \frac{v+h}{\sqrt{2}}\end{pmatrix}.$ (3)
Once Higgs gets a vacuum expectation value, $v$, it induces mixing between singlet and neutral component of doublet fermion through $\overline{\Psi}\tilde{H}\chi$-coupling. Denoting the mass eigenstates as $\chi_{0}$ and $\chi_{1}$ with the mixing angle $\sin\theta$, the transformation from the flavor states to physical states can be written as,
$\left(\begin{matrix}\psi^{0}\\ \chi\end{matrix}\right)=\left(\begin{matrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{matrix}\right)\left(\begin{matrix}\chi_{0}\\ \chi_{1}\end{matrix}\right),$ (4)
where the mixing angle is given by
$\tan{2\theta}=\frac{\sqrt{2}yv}{M_{\Psi}-M_{\chi}}.$ (5)
The mass eigenvalues of the physical states are given as
$M_{\chi_{0}}$ $=$ $M_{\Psi}\cos^{2}\theta+\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\sin^{2}\theta,~{}~{}$
$M_{\chi_{1}}$ $=$ $M_{\Psi}\sin^{2}\theta-\frac{yv}{\sqrt{2}}\sin 2\theta+M_{\chi}\cos^{2}\theta\equiv M_{\rm DM}$ (6)
with a mass-splitting $\Delta M=M_{\chi_{0}}-M_{\chi_{1}}$, where, $\chi_{1}$ is dominantly the singlet fermion and $\chi_{0}$ is the $\psi^{0}$. From Eq. (1), the interaction among the mass eigenstates ($\chi_{0}$, $\psi^{\pm}$ and $\chi_{1}$), can be expressed as,
$\mathcal{L}_{int}$ $=$ $\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\left[\sin^{2}\theta\overline{\chi_{1}}\gamma^{\mu}Z_{\mu}\chi_{1}+\cos^{2}\theta\overline{\chi_{0}}\gamma^{\mu}Z_{\mu}\chi_{0}\right.$ (7)
$\left.+\sin\theta\cos\theta(\overline{\chi_{1}}\gamma^{\mu}Z_{\mu}\chi_{0}+\overline{\chi_{0}}\gamma^{\mu}Z_{\mu}\chi_{1})\right]$
$+\frac{e_{0}}{\sqrt{2}\sin\theta_{W}}\sin\theta\overline{\chi_{1}}\gamma^{\mu}W_{\mu}^{+}\psi^{-}+\frac{e_{0}}{\sqrt{2}\sin\theta_{W}}\cos\theta\overline{\chi_{0}}\gamma^{\mu}W_{\mu}^{+}\psi^{-}$
$+\frac{e_{0}}{\sqrt{2}\sin\theta_{W}}\sin\theta{\psi^{+}}\gamma^{\mu}W_{\mu}^{-}\chi_{1}+\frac{e_{0}}{\sqrt{2}\sin\theta_{W}}\cos\theta{\chi^{+}}\gamma^{\mu}W_{\mu}^{-}\chi_{0}$
$-\left(\frac{e_{0}}{2\sin\theta_{W}\cos\theta_{W}}\right)\cos 2\theta_{W}{\psi^{+}}\gamma^{\mu}Z_{\mu}\chi^{-}-e_{0}{\psi^{+}}\gamma^{\mu}A_{\mu}\psi^{-}$
$-\frac{y}{\sqrt{2}}h\left[\sin 2\theta(\overline{\chi_{1}}\chi_{1}-\overline{\chi_{0}}\chi_{0})+\cos 2\theta(\overline{\chi_{1}}\chi_{0}+\overline{\chi_{0}}\chi_{1})\right],$
where $e_{0}=0.313$ is the electromagnetic coupling constant, and $\sin\theta_{W}$ is the Weinberg angle.
3 Dark matter phenomenology
In this Section, we systematically explore the relic density parameter space for SDDM, considering annihilation, co-annihilation, co-scattering, and freeze-in, depending on the $\sin\theta$ and $\Delta{M}$. We divide the parameter space into three regimes: (i) thermal DM relic considering annihilation and co-annihilation only, (ii) thermal DM relic via co-annihilation and co-scattering processes, and (iii) freeze-in production of the DM.
3.1 Thermal dark matter relic considering annihilation and co-annihilation only
Here, we assume that the DM is in thermal equilibrium with the SM bath particles. This can be realized by considering a large SD mixing. The SD mixing is chosen in such a way that all the dark sector particles decouple at the same epoch. The relic density of DM in this scenario can be achieved by solving the Boltzmann equation (BE) [39],
${}\frac{dn}{dt}+3\mathcal{H}n=-\langle\sigma v\rangle_{\rm eff}(n^{2}-(n^{\rm eq})^{2})$ (8)
where $n=\sum_{i}n_{i}$ represents the total number density of all the dark sector particles, $n^{\rm eq}$ is the equilibrium number density, and the Hubble parameter is expressed as $\mathcal{H}=\sqrt{\frac{4\pi^{2}g_{*\rho}}{45M_{Pl}}}T^{2}$ with the Planck mass, $M_{pl}=1.22\times 10^{19}$ GeV, and $g_{*\rho}$ is the total relativistic d.o.f contributing to the energy density. $\langle\sigma v\rangle_{\rm eff}$ represents the effective cross-section which takes into account all number changing processes for DM freeze-out is given by
$\langle\sigma v\rangle_{\rm eff}$ $=$ $\frac{g^{2}_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{1}}\chi_{{}_{1}}}+\frac{g_{0}g_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{0}}\chi_{{}_{1}}}(1+\Delta_{\chi_{{}_{0}}})^{{3}/{2}}\exp(-x\Delta_{\chi_{{}_{0}}})$ (9)
$+\frac{g_{2}g_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{1}}\psi^{-}}(1+\Delta_{\psi^{-}})^{{3}/{2}}\exp(-x\Delta_{\psi^{-}})$
$+\frac{g^{2}_{1}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{0}}\chi_{{}_{0}}}(1+\Delta_{\chi_{{}_{0}}})^{3}\exp(-2x\Delta_{\chi_{{}_{0}}})+\frac{g^{2}_{2}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\psi^{+}\psi^{-}}(1+\Delta_{\psi^{-}})^{3}\exp(-2x\Delta_{\psi^{-}})$
$+\frac{g_{0}g_{2}}{g^{2}_{\rm eff}}\langle\sigma v\rangle_{\chi_{{}_{0}}\psi^{-}}(1+\Delta_{\chi_{{}_{0}}})^{3/2}(1+\Delta_{\psi^{-}})^{3/2}\exp(-x(\Delta_{\chi_{{}_{0}}}+\Delta_{\psi^{-}}))$
where $g_{0},g_{1}$ and $g_{2}$ represent the internal d.o.f. of $\chi_{0},\chi_{1}$ and $\psi^{-}$ respectively and $\Delta_{i}$ stands for the ratio $(M_{i}-M_{\chi_{1}})/M_{\chi_{1}}$ with $M_{i}$ denoting the mass of $\chi_{0}$ and $\psi^{-}$. Here $g_{\rm eff}$ is the effective degree of freedom which is given by:
$g_{\rm eff}=g_{1}+g_{0}(1+\Delta_{\chi_{{}_{0}}})^{3/2}\exp(-x\Delta_{\chi_{{}_{0}}})+g_{2}(1+\Delta_{\psi^{-}})^{3/2}\exp(-x\Delta_{\psi^{-}})$ (10)
and $x$ is the dimensionless parameter $M_{\rm\chi_{1}}/T$. The relic density of DM $\chi_{1}$ can then be obtained as
[…]
Here $x_{f}=M_{\chi_{1}}/T_{f}$, and $T_{f}$ denotes the freeze-out temperature of $\chi_{1}$.
Upon analyzing Eq. (9), the first three terms correspond to the annihilation and co-annihilation of $\chi_{1}$, while the last three terms describe the annihilation and co-annihilation processes involving $\chi_{0}$ and $\psi^{-}$. The latter terms are associated with gauge-mediated interactions and are independent of the mixing angle: $\sin\theta$. Consequently, these processes dominate over the first three terms in the regime of smaller values of $\sin\theta$. Therefore, it is essential to verify whether the DM remains in chemical equilibrium with the $\mathcal{Z}_{2}$-odd doublet fermion in this low $\sin\theta$ region. We define a new parameter, $\rho_{1}$, to validate the assumption of chemical equilibrium among the dark sector particles,
$\rho_{1}=\frac{\Omega_{1s}h^{2}}{\Omega_{2s}h^{2}(\rm no~{}coscattering)},$ (12)
where $\Omega_{1s}h^{2}$ is the relic density obtained by solving Eq (8), and $\Omega_{2s}h^{2}$(no co-scattering) is calculated by solving Eqs (LABEL:eq:Y1) and (LABEL:eq:Y2) while excluding co-scattering processes [^1] We use micrOMEGAs 6.1.15[40, 35] to solve the Boltzmann equations, where darkOmega function is used to estimate $\Omega_{1s}h^{2}$ and darkOmegaN is used to estimate $\Omega_{2s}h^{2}$(no co-scattering) while excluding the co-scattering processes.. In Eq. (12), $\rho_{1}=1$ implies all the dark sector particles decouple at the same epoch. In other words, solving Eq. (8) or Eqs. (LABEL:eq:Y1) and (LABEL:eq:Y2) are equivalent. $\rho_{1}<1$ implies the dark sector particles are decoupling at different epochs. When $\sin\theta$ is small, the singlet decouples earlier than the doublet and produces a relatively larger relic. Therefore, one expects $\rho_{1}\leq 1$.
Figure 1: Allowed parameter space of DM for which DM can be considered to be in equilibrium with SM bath and the $\mathcal{Z}_{2}$-odd doublet. All the points in the plots satisfy the required relic density. The grey-shaded region in the plot shows the LEP bound [41] on the heavy fermion doublet.
We have presented the parameter space in Fig. 1. In this plot, all the points satisfy the correct relic density of DM, and we have used $0.99<\rho_{1}\leq 1$ to assess the validity and domain of applicability of the Eq. (8). In Fig. 1 (left), we have shown all the points in the plane of $\sin\theta$ and $M_{\rm DM}$, and in the right plot, we projected the same points in the plane of $\Delta{M}$ and $M_{\rm DM}$. DM mass is varied in the range of $1-1000$ GeV, $\Delta{M}$ is varied in the range of $1-1000$ GeV, and $\sin\theta$ is varied in the range of ($0.5-10^{-4}$). From Fig. 1 (left), we observe that the value of $\sin\theta$ can be as small as $\sim 5\times 10^{-4}$ for $M_{\rm DM}\sim 250$ GeV. Note that all these points in Fig. 1 correspond to the simultaneous decoupling of singlet and doublet from the thermal bath. As shown in Fig. 1 (right), we exactly reproduced all the points that satisfy correct relic density in the plane of $\Delta{M}$ vs $M_{\rm DM}$ [8].
3.2 Thermal dark matter relic via co-annihilation and
co-scattering processes
We have already demonstrated in Fig. 1 (left) the parameter space for the correct relic of DM in $\sin\theta$-$M_{\rm DM}$ plane, where the singlet and the doublet freeze-out at the same epoch through annihilation and co-annihilation processes. Beyond this regime, the singlet may reach equilibrium. However, in this case, the singlet and doublet will decouple at different epochs. Due to small mixing, the singlet will decouple early, while the doublet components will remain in equilibrium for a longer period due to their gauge interactions. Even though the singlet decouples early, the freeze-out abundance of DM may alter because of the presence of additional processes, such as co-scattering or conversion-driven freeze-out [32, 33], which were subdominant for larger SD mixing. As a result, different sets of evolution equations for the dark sector particles have to be solved to get the relic density. In the SDDM model, we define two dark sectors: (a) sector 1, which contains $\chi_{1}$, and (b) sector 2, which contains $\chi_{0},\psi^{-}$, whereas all the SM particles are assigned as sector 0. The singlet fermion, being the lightest among the dark sector particles, serves as the DM candidate. By defining the comoving number density of sector 1 and sector 2 particles as $Y_{1}$ and $Y_{2}$  [^2] Processes like $\psi^{-}SM\rightarrow\chi_{0}SM$ will keep the neutral and charged components of the doublet in equilibrium always which makes the abundances of sector 2 particles to be the same., respectively, the general coupled BEs for their evolution are given as
$\frac{dY_{1}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{1100}v\rangle(Y_{1}^{2}-{Y_{1}^{eq}}^{2})+\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})\right.$
$+\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{eq}}{Y_{2}^{eq}}\right)-\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)-\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right],$
$\frac{dY_{2}}{dT}$ $=$ $\frac{1}{3\mathcal{H}}\frac{ds}{dT}\left[\langle\sigma_{2200}v\rangle(Y_{2}^{2}-{Y_{2}^{eq}}^{2})-\langle\sigma_{1122}v\rangle\left(Y_{1}^{2}-Y_{2}^{2}\frac{{Y_{1}^{eq}}^{2}}{{Y_{2}^{eq}}^{2}}\right)+\langle\sigma_{1200}v\rangle(Y_{1}Y_{2}-Y_{1}^{eq}Y_{2}^{eq})\right.$
$-\left.\langle\sigma_{1222}v\rangle\left(Y_{1}Y_{2}-Y_{2}^{2}\frac{Y_{1}^{eq}}{Y_{2}^{eq}}\right)+\langle\sigma_{1211}v\rangle\left(Y_{1}Y_{2}-Y_{1}^{2}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right],$
where $Y_{i}^{eq}$ are the equilibrium abundances, $\mathcal{H}$ is the Hubble parameter, $\langle\sigma_{\alpha\beta\gamma\delta}v\rangle$ are the thermally averaged cross-sections for processes involving annihilation of particles of sectors $\alpha\beta\rightarrow\gamma\delta$. $\Gamma_{2\rightarrow 1}$ is the conversion term, which includes both the interaction rate of the co-scattering process ($\langle\sigma_{1020}v\rangle$) and the decay ($\Gamma_{D}$).
In the large mixing angle regime, only annihilation and co-annihilations are the dominating processes that decide the relic. In this case, the decoupling of the particles in both sectors happens in the same epoch, and these two equations simplify to the one-sector BE as given in Eq.(8). On the other hand, in the small SD mixing regime, annihilation and co-annihilation processes are sub-dominant. As a result Eq. (LABEL:eq:Y1) and (LABEL:eq:Y2) are simplified to
$\frac{dY_{1}}{dT}=-\frac{\Gamma_{2\rightarrow 1}}{\mathcal{H}T}\left[Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right],$ (15)
$\frac{dY_{2}}{dT}=\frac{s}{\mathcal{H}T}\left[\left<\sigma_{2200}v\right>\left(Y^{2}_{2}-{Y^{eq}_{2}}^{2}\right)+\frac{\Gamma_{2\rightarrow 1}}{s}\left(Y_{2}-Y_{1}\frac{Y_{2}^{eq}}{Y_{1}^{eq}}\right)\right].$ (16)
We have shown the evolution of the abundances of both sectors w.r.t. temperature for two different mixing angles: $\sin\theta=1.19\times 10^{-1}$ and $\sin\theta=1.19\times 10^{-5}$ in Fig. 2. We fixed the DM mass to 146.48 GeV and took the mass splitting to be 10.5 GeV and 8.84 GeV, respectively, to get the correct relic abundances in both cases. For the large mixing angle, we see that both sectors remain in equilibrium for around $T\sim 6$ GeV in the left panel of Fig. 2. In this case, co-annihilation is the dominant process, which is deciding the relic. When we decrease the mixing angle by four orders of magnitudes, the co-annihilation and annihilation processes, which depend on the mixing angle as $\sin^{2}\theta$ and $\sin^{4}\theta$ respectively, become negligible. We see that the sector 1 particle goes out of equilibrium very early at around $T\sim 8$ GeV, whereas the sector 2 particles are still in equilibrium due to their gauge interactions with the SM bath in right panel of Fig. 2. The sector 1 particle goes out of equilibrium with a larger abundance, then upscatters with the SM particles to the sector 2 particles, which is the co-scattering process. Thus, the relic of sector 1 comes down to the correct ballpark.
Figure 2: Evolutions of sector 1 and sector 2 particle abundances for $M_{\rm DM}$=146.48 (146.48) GeV, $M_{\psi}=156.98~{}(155.32)$ GeV, and $\sin\theta=1.19\times 10^{-1}~{}(1.19\times 10^{-5})$ are shown in left (right) plot which gives the DM relic of $\Omega_{\rm DM}h^{2}\simeq 0.12$.
We now explore the regime of the co-scattering effects in the SDDM scenario, which was not taken into account in earlier works [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18]. To explore how the relic density of DM depends on our model parameters, we performed several analyses and examined the allowed parameter space.
Figure 3: Variation of DM relic density w.r.t. DM mass with the fixed mass splitting of $51<\Delta{M}<$ 5 GeV for $10^{-2}<\sin\theta<10^{-1}$ [left] and $10^{-5}<\sin\theta<10^{-4}$ [right] considering co-scattering processes (red) and without considering co-scattering processes (blue).
We study the relic density of DM considering co-scattering and without co-scattering with different sets of our model parameters. In Fig. 3, we have kept the mass splitting in the range of 1 GeV to 5 GeV and then varied the mixing angle in two sets. In the left pannel of Fig. 3, the mixing angle is varied in between $10^{-2}-10^{-1}$. Due to smaller mass splitting and the large mixing angle, the co-annihilations are the most dominating processes that determine the relic of the DM. Co-scattering effects are sub-dominant in this region as the mixing angle is large enough to keep the DM in the thermal equilibrium for a longer period. The freeze-out via co-annihilation processes finally decides the relic of the DM in this case. We see that the relic density points obtained considering the co-scattering and without co-scattering coincide, validating the co-annihilation regime. Now, keeping the mass splitting in the same range, we decrease the mixing angle to a range of $10^{-5}-10^{-4}$. Now, because of this small mixing angle, the annihilation/co-annihilation processes become subdominant, and the DM gets chemically decoupled, leading to a larger DM abundance. These points are shown with the blue dots in the right panel of Fig. 3. Now, when we consider the co-scattering effects, the dynamics of the freeze-out of DM change drastically. The DM can now scatter with SM particles to a higher mass state, $\psi^{\pm},\chi_{0}$, and SM particles, known as co-scattering processes or conversion-driven processes, followed by annihilation of doublets to SM particles. These processes help the DM to bring down its relic density to the correct ballpark. These points are shown in red color in the right panel of Fig. 3. We see that the co-scattering effects are significant in the small mixing angle regime while they are sub-dominant for the large mixing angles.
Figure 4: Variation of DM relic density w.r.t DM mass with the fixed mass splitting of $105<\Delta{M}<$ 10 GeV for $10^{-2}<\sin\theta<10^{-1}$ [left] and $10^{-5}<\sin\theta<10^{-4}$ [right] considering co-scattering processes (red) and without considering co-scattering processes (blue).
We now move to Fig. 4, where we have varied the mass splitting in the range from 5 to 10 GeV, while the SD mixing is kept in the same range as in Fig. 3. We observe the dependency of the DM relic on the SD mixing angles remains similar as in Fig. 3. When we increase the mass splitting, the co-annihilation or co-scattering cross becomes smaller, which leads to larger DM abundance. This is clearly visible from the left and right panel of Fig. 4.
Figure 5: Correct relic parameter space in the plane of $\Delta{M}-M_{\rm DM}$ excluding co-scattering (left) and including co-scattering (right) with different ranges of mixing angles as mentioned in the plot.
In Fig. 5, we present the points that yield the correct DM relic density in the $\Delta M$-$M_{\rm DM}$ plane. In the left plot, we show the points that give the correct relic density when co-scattering processes are excluded. The relic density calculation considers only annihilation, co-annihilation, and decay processes. The color coding represents different values of $\sin\theta$, ranging from 0.5 to $5\times 10^{-4}$, as indicated in the plot. As observed, for $\sin\theta>0.05$, the results are consistent with previous findings [8]. For $\sin\theta\leq 0.05$, we find new parameter space shown in magenta and cyan color points. For small $\Delta M$, decay and inverse-decay processes are comparable to co-annihilation processes. For a typical SD mixing, represented by magenta-colored points, the cross-section decreases as we move from left to right due to the increasing DM mass. This decrease can be compensated by reducing $\Delta M$. For larger $\Delta M$ [e.g., $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$], the co-annihilation cross-section experiences greater suppression, causing DM to decouple earlier and resulting in a larger abundance. However, with decay terms ($\Gamma_{D}$) included, the DM relic density can be brought into the correct range. Table 1 shows that for $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$, the DM is overabundant in the absence of decay. In the presence of decay and inverse-decay processes in the Boltzmann equation, DM remains in equilibrium for a longer duration, reducing the relic abundance to the desired value. The evolution of yields both considering decay and without decay has been shown in Fig. 12 in Appendix A. In the absence of decay terms, a small $\Delta M$ [e.g., $B({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\star})$] delays the DM decoupling epoch, resulting in a smaller yield ($Y_{1}(B)<Y_{1}(A)$). In this scenario, a smaller contribution from decay and inverse-decay processes is sufficient to achieve the correct relic density. This is evident in Table 1, where turning off the decay term results in a smaller increase in relic density compared to $A({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\star})$. For even smaller $\Delta M$ [e.g., $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$], co-annihilation combined with decay-inverse decay fails to maintain equilibrium, leading to DM freezing out with a larger abundance. This behavior can be understood from $\Gamma_{D}\propto\sin^{2}\theta\Delta M^{5}$. For a fixed $\sin\theta$, as at point $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$, $\Delta M$ is too small to sustain equilibrium. Increasing $\sin\theta$ allows $\Gamma_{D}$ to adjust the relic density, as seen for $C({\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\star})$ with $\sin\theta=9.9\times 10^{-2}$. The magenta-colored points curve back leftward from $B({\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\star})$. In this regime, decreasing DM mass and $\Delta M$ increases the cross-section, leading to a late decoupling of DM with a comparably smaller abundance. In this case, the contributions from decay terms are sufficient enough to provide the correct relic density. The turning point varies with $\sin\theta$: for smaller SD mixing, it occurs at smaller DM masses, while for larger SD mixing, it occurs at larger DM masses.
Point $M_{\rm DM}$ (GeV) $\Delta{M}$ (GeV) $\sin\theta$ $\Omega_{2s}h^{2}$(no co-scattering) $\Omega_{2s}h^{2}$(no co-scattering) without decay
[…]
Table 1: Relic density for benchmark points: (a) without considering co-scattering, and (b) without considering co-scattering and decay.
As discussed earlier, the co-scattering is important for small values of $\sin\theta$. In the right panel of Fig. 5, we present the correct relic density points, including the contribution from co-scattering processes. The parameter space that previously yielded the correct relic density when co-scattering was excluded gets modified due to additional processes. Conversely, the over-abundant points with $\sin\theta<10^{-4}$, which did not provide the correct relic density in the absence of co-scattering processes, now yield the correct relic density once these processes are included. For further clarification, we define a new parameter,
$\rho_{2}=\frac{\Omega_{2s}h^{2}}{\Omega_{2s}h^{2}({\rm no~{}coscattering})},$ (17)
where $\Omega_{2s}h^{2}$ represents relic density, including co-scattering, while $\Omega_{2s}h^{2}$(no co-scattering) denotes relic density in the absence of the co-scattering contribution. From the definition of $\rho_{2}$ given in Eq. (17), $\rho_{2}=1$ indicates the relic density is independent of co-scattering effects. Consequently, any deviation from 1 will indicate the role of co-scattering in the final relic density. The larger is the deviation, the more is the effect from co-scattering. In fact, $\Omega_{2s}h^{2}$ includes all the processes involving annihilation, co-annihilation, co-scattering, and decay, while $\Omega_{2s}h^{2}$ (no co-scattering) involves only annihilation, co-annihilation, and decay. This implies the effective cross-section, giving rise to $\Omega_{2s}h^{2}$, is always larger than or equal to $\Omega_{2s}h^{2}$ (no co-scattering). Therefore, $\Omega_{2s}h^{2}\leq\Omega_{2s}h^{2}$ (no co-scattering), resulting in $\rho_{2}\leq 1$.
Figure 6: Co-scattering parameter space in the plane of $\sin\theta$ vs $M_{\rm DM}$. The color code represents the parameter $\rho_{2}$.
In Fig. 6, we present all the points that satisfy correct relic density in the plane of $\sin\theta$ vs $M_{\rm DM}$. The range of $\sin\theta$ is chosen in such a way that the DM will be in equilibrium. The color code depicts the $\rho_{2}$ values. If $\rho_{2}$=1, then co-scattering is sub-dominant, which is shown by the violet-colored points. As we decrease the value of $\sin\theta$, the $\rho_{2}$ deviates from 1, which implies the nontrivial role of co-scattering in large parameter space shown by cyan, yellow, and purple color points.
3.3 Freeze-in production of dark matter
In Sections 3.1, and 3.2, we saw that the singlet is brought to equilibrium for the $\sin\theta\gtrsim\mathcal{O}(10^{-6})$. This implies that for $\sin\theta\gtrsim\mathcal{O}(10^{-6})$, the relic density of SDDM will be obtained via the freeze-out mechanism. If the SD mixing is further reduced, the DM relic density can no longer be determined through the conventional freeze-out mechanisms. However, it can be estimated through the freeze-in mechanism. In this regime, the heavier $\mathcal{Z}_{2}$-odd particles (belonging to sector 2) decay into DM (belonging to sector 1) in equilibrium, causing the DM number density to be produced via the freeze-in mechanism. The equation for the freeze-in mechanism, where particles from sector 2 decay into DM, is given by
$\frac{dY_{1}}{dx}=\frac{\Gamma_{tot}}{\mathcal{H}x}\frac{K_{1}(x)}{K_{2}(x)}Y^{eq}_{2},$ (18)
where $x=M_{\rm DM}/T$, and $\Gamma_{tot}$ is the total decay rate, which includes the two body decays like $\Gamma_{\chi_{0}\rightarrow\chi_{1}H},\Gamma_{\chi_{0}\rightarrow\chi_{1}Z},\Gamma_{\psi^{\pm}\rightarrow\chi_{1}W^{\pm}}$, and three body decay of the doublets as $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}l^{\pm}\nu_{l}}$ [^3] The decay mode $\Gamma_{\psi^{\pm}\rightarrow\chi_{1}\pi^{\pm}}$ is very suppressed. This is because of $\Delta M\geq 1$ GeV.. The decay rates are given in Appendix B.
Figure 7: Parameter space for the correct relic density through freeze-in processes in the plane of $\sin\theta$ vs DM mass. The color code represents the mass splitting between the DM and next to the lightest $\mathcal{Z}_{2}$ odd particle.
In Fig. 7, we have shown the correct relic density points in the plane of $\sin\theta$ vs $M_{\rm DM}$ where the color code represents $\Delta M$. When $\Delta M\lesssim 80$ GeV, the freeze-in relic density is primarily determined by three-body decay modes. Conversely, as $\Delta M$ exceeds 80 GeV, 91 GeV, and 125 GeV, the two-body decay channels $\psi^{\pm}\rightarrow\chi_{1}W^{\pm}$, $\chi_{0}\rightarrow\chi_{1}Z$, and $\chi_{0}\rightarrow\chi_{1}H$ become accessible, respectively.
The three-body decay width is $\propto\sin^{2}\theta\Delta M^{5}$. When the $\Delta M$ increases, $\sin\theta$ has to be decreased to give the required decay width to produce the correct relic density. This trend is evident in Fig. 7. When the two body decay modes dominate in the larger mass-splitting regime, the decay width increases with an increase in $M_{\rm DM}$, $\Delta M$, and $\sin\theta$. For a fixed $\Delta M$, an increase in $M_{\rm DM}$ requires a reduction in SD mixing to achieve the correct relic density. Similarly, for a fixed $M_{\rm DM}$, an increase in $\Delta M$ necessitates adjusting the SD mixing to obtain the decay width that ensures the correct relic density. This behavior is illustrated in Fig. 7.
4 Direct detection of dark matter
The DM can scatter off the target nucleus via Higgs and Z boson mediated processes as shown in Fig. 8. This leads to the spin-independent (SI) cross-section for Higgs mediated dark matter-nucleon scattering to be
$\sigma_{h}^{\rm SI}$ $=\frac{4}{\pi A^{2}}\mu^{2}_{r}\frac{\Delta M^{2}\sin^{4}2\theta}{2v^{2}M^{4}_{h}}\Big{[}\frac{m_{p}}{v}\Big{(}f^{p}_{Tu}+f^{p}_{Td}+f^{p}_{Ts}+\frac{2}{9}f^{p}_{TG}+\frac{m_{n}}{v}\Big{(}f^{n}_{Tu}+f^{n}_{Td}+f^{n}_{Ts}+\frac{2}{9}f^{n}_{TG}\Big{)}\Big{]}^{2}.$ (19)
Different coupling strengths between DM and light quarks are given by [42, 43] as $f^{p}_{Tu}=0.020\pm 0.004,~{}f^{p}_{Td}=0.026\pm 0.005,~{}f^{p}_{Ts}=0.014\pm 0.062,~{}f^{n}_{Tu}=0.020\pm 0.004,~{}f^{n}_{Td}=0.036\pm 0.005,~{}f^{n}_{Ts}=0.118\pm 0.062$. The coupling of DM with the gluons in target nuclei are parameterized by [44] $f^{p,n}_{TG}=1-\sum_{q=u,d,s}f^{p,n}_{Tq}$. Similarly, the SI DM-nucleon scattering cross-section through Z mediation is given by [45, 46]:
$\sigma^{\rm SI}_{Z}=\frac{G^{2}_{F}\sin^{4}\theta}{\pi A^{2}}\mu_{r}^{2}\Big{|}\left[Zf_{p}+(A-Z)f_{n}\right]^{2}\Big{|}^{2},$ (20)
where the $f_{p}=f_{n}=0.33$ corresponds to the form factors for proton and neutron, respectively. Here $\mu_{r}$ is the reduced mass of the DM-nucleon system, and $A$ and $Z$ are the mass number and atomic number, respectively.
Figure 8: Feynman diagram for spin-independent DM-nucleon scattering cross-section.
A larger mixing angle leads to an increased direct detection cross-section, which is excluded by the direct detection experiment like LZ [6]. We calculated the SI direct detection cross-section ($\sigma^{\rm SI}_{\rm DM-N}=\sigma_{h}^{\rm SI}+\sigma^{\rm SI}_{Z}$) for the correct relic points and shown w.r.t $M_{\rm DM}$ in Fig. 9. The direct detection excludes the SD mixing angle for $\sin\theta>0.07$. We have also shown the LZ exclusion region in the parameter space of $\sin\theta$ vs $M_{\rm DM}$ in Fig. 11. The DARWIN [47] sensitivity is shown with the blue dashed line. When we include the LEP bound, the exclusion goes to $\sin\theta>0.04$.
Figure 9: Spin-independent DM-nucleon scattering cross-section as a function of DM mass.
5 Displaced vertex signatures
Because of gauge interactions, the doublet components, $\psi^{\pm}$, and $\chi_{0}$ can be copiously produced in the colliders. Once these particles are produced, they will decay to the DM and charged lepton via the off-shell gauge bosons in the small mass splitting range after traveling some finite distance. The decay length is $\propto(\sin^{2}\theta\Delta{M}^{5})^{-1}$. Thus, one can measure the displaced vertex signatures in present and future collider experiments such as LHC and MATHUSLA. The inclusion of co-scattering can enhance the reach of these experiments in searching for these dark sector particles. We have calculated the decay lengths of such particles considering mass splitting ($\Delta M$) in the range from 1 to 12 GeV and shown in Fig. 10 with varying DM mass. We considered three ranges of mixing angles: $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$, $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$, and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$. The three patches correspond to these mixing angles in the bottom to top order, respectively. The color code represents the parameter $\rho_{2}$. The figure depicts that the inclusion of the co-scattering enhances the lifetime of the dark partners and hence reaches the decay length of LHC and MATHUSLA.
Figure 10: Decay length of the doublet fermion as a function of DM mass with $\rho_{2}$ in the color code. The mixing angles are varied in the range of $10^{-2}\leq\sin\theta\leq 2\times 10^{-2}$, $10^{-4}\leq\sin\theta\leq 2\times 10^{-4}$, and $10^{-5}\leq\sin\theta\leq 2\times 10^{-5}$ and the three patches correspond to these mixing angles in the bottom to top order, respectively.
6 Conclusions
In this work, we explored the full parameter space of singlet-doublet vector-like fermionic dark matter for the first time, accounting for annihilation, co-annihilation, co-scattering, and freeze-in processes. In Fig. 11, we show the parameter space in the plane of mixing angle versus DM mass, where the black dashed line separates the thermal and non-thermal regimes at $\Delta M$ = 1 GeV. For $\Delta M<$ 1 GeV, this line shifts downward, and for $\Delta M>$ 1 GeV, this line moves upward. In the upper part of the line, the DM can attain thermal equilibrium, and the relic can be obtained via freeze-out mechanism. On the other hand, below this black dashed line, the DM can never reach equilibrium, and the relic can be obtained via the freeze-in mechanism only. The grey-shaded region represents the parameter space where the annihilation and co-annihilation processes solely decide the relic. When the mixing angle drops further, the co-scattering processes become the most dominant one, and they decide the DM relic in this regime. This is shown with the green-shaded region.
Figure 11: Phase diagram of the SDDM parameter space. The black dashed line separates the region of thermal (upper) and non-thermal (lower) regimes.
As $\sin\theta$ decreases further, the DM decouples from equilibrium earlier, freezing out with a larger abundance. This over-abundant region is depicted as the cyan-shaded area. When the mixing angle drops below $<2\times 10^{-7}$, thermal production of DM becomes inefficient. Below the black dashed line, the doublet decays into the singlet DM, which subsequently freezes into the correct relic abundance. This freeze-in regime is shown in the blue-shaded region. However, for $\sin\theta<\mathcal{O}(10^{-13})$, the doublet components fail to produce enough singlet DM, resulting in an under-produced region shown in white.
Direct detection experiments constrain the SD mixing angle to values below 0.07, and the LEP bounds on the doublet fermion further tighten this exclusion to 0.04. We have also investigated the displaced vertex signatures of our model at the LHC and MATHUSLA. Our analysis shows that including co-scattering processes significantly expands the parameter space accessible to both LHC and MATHUSLA experiments.
7 Acknowledgments
P.K.P. would like to acknowledge the Ministry of Education, Government of India, for providing financial support for his research via the Prime Minister’s Research Fellowship (PMRF) scheme. The work of N.S. is supported by the Department of Atomic Energy-Board of Research in Nuclear Sciences, Government of India (Ref. Number: 58/14/15/2021- BRNS/37220). We would like to thank Alexander Pukhov for his assistance on $\tt micrOMEGAs$.
Appendix A Evolution of abundances for point A
[…]
Appendix B Decay rates
Figure 13: The Feynman diagrams for the production of the DM from the decay of sector 2 particles.
Two body decay rates are given as
$\Gamma_{a\rightarrow bc}$ $=$ $\frac{\sqrt{\Lambda(M_{a}^{2},M_{b}^{2},M_{c}^{2})}}{2M_{a}}\frac{1}{32\pi^{2}M_{a}^{2}}4\pi\left|\mathcal{M}_{a\rightarrow bc}\right|^{2},$ (21)
${\rm where,~{}}\Lambda(x,y,z)$ $=$ $x^{2}+y^{2}+z^{2}-2xy-2yz-2xz$ (22)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi H}\right|^{2}$ $=$ $\frac{(M_{\chi_{0}}-M_{\chi_{1}})^{2}\sin^{2}2\theta}{4v^{2}}\left(1-2\sin^{2}\theta\right)^{2}\left(M_{\chi_{1}}^{2}+2M_{\chi_{0}}M_{\chi_{1}}+M_{\chi_{0}}^{2}-M_{h}^{2}\right),$ (23)
$\left|\mathcal{M}_{\psi^{0}\rightarrow\chi Z}\right|^{2}$ $=$ $G_{F}\left(1-\sin^{2}\theta\right)\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\chi_{0}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{Z}-6M_{\chi_{0}}M_{\chi_{1}}M^{2}_{Z}+M^{4}_{\chi_{0}}$ (24)
$+M^{2}_{\chi_{0}}M^{2}_{Z}-2M^{4}_{Z}),$
$\left|\mathcal{M}_{\psi^{-}\rightarrow\chi W}\right|^{2}$ $=$ $\frac{G_{F}}{2}\sin^{2}\theta(M^{4}_{\chi_{1}}-2M^{2}_{\psi^{-}}M^{2}_{\chi_{1}}+M^{2}_{\chi_{1}}M^{2}_{W}-6M_{\psi^{-}}M_{\chi_{1}}M^{2}_{W}+M^{4}_{\psi^{-}}$ (25)
$+M^{2}_{\psi^{-}}M^{2}_{Z}-2M^{4}_{Z}).$
The three body decay rates are [48]
$\Gamma_{\psi\rightarrow\chi l\nu_{l}}=\frac{2G_{F}^{2}}{15\pi^{3}}(\Delta M)^{5}$ (26)